Jami, Ryan, and Sandra decided to create a riddle to reveal their ages. Jami is 5 years older than Ryan. Ryan is 7 years older than Sandra. The sum of their ages is 73. Write equations for this situation and use the equations to determine each person’s age.

Found 2 solutions by mananth, josgarithmetic:
Answer by mananth(16949)


Jami, Ryan, and Sandra decided to create a riddle to reveal their ages. Jami is 5 years older than Ryan. Ryan is 7 years older than Sandra. The sum of their ages is 73. Write equations for this situation and use the equations to determine each person’s age.
let sandra be x years
Ryan age = x+7
Jami,s age =x+12
x+x+7+x+12=73
3x+19=73
3x = 54
x= 18 Sandra's age


Answer by josgarithmetic(39838)

Jami              x+7+5
Ryan              x+7
Sandra             x
SUM              73

.
.
.